Diffusion Morning Report 02/28/05

Ted Berg tberg at snf.stanford.edu
Mon Feb 28 06:46:54 PST 2005


Good Morning All,
    Looking  good this morning.  LTO had a possible leak issue-looking 
into it.  Epi had a bad lamp-replaced. TEOS2 band Thermco poly testing.
Everything else looks OK for now.

  	Date: 	2/28/2005 	THICKNESS-Change at
Tube Name/Number 	Status 	Comments 	 
Tylan 1 	Up 	  	 
Tylan 2 	Up 	  	 
Tylan 3 	Up 	  	 
Tylan 4 	Up 	  	 
Tylan 5-BBr3 	Up 	  	 
Tylan 6-POCl3 	Up 	  	 
Tylan 7-TEOS2 	down 	Qual testing 	 
Tylan 8-SiGe 	Up 	  	1.41/30
Tylan 9-Poly 	Up 	  	22.65/30
Tylan 10-Nitride 	Up 	  	22.77/30
Tylan 11-BPSG-LTO 	Up 	pull and clean completed 	1.25/30
Tylan 12-FGA 	Up 	  	 
Tystar 	yellow 	  	31.61/30
Doped Poly 	GONE 	  	 
TEOS 	yellow 	  	 
Tylan 16 	N/A 	  	 
Thermco 1 	Up 	  	 
Thermco 2 	Up 	  	 
Thermco Poly 	yellow 	testing doped 	 
Thermco Nitride 	Down 	Waiting 	 
  	  	  	 
Burn Box 	Up 	  	 
ZYGO 	up 	  	 
  	  	  	 
GAAS 22 	UP 	  	 
GAAS 23 	UP 	  	 
  	  	  	 
SiH4 for furnaces 	  	Auto changeover working well  	A on line 180 PSI


-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20050228/52c1b9b9/attachment.html>


More information about the diffusion mailing list